# Augmenting Molecular Graphs with Geometries via Machine Learning Interatomic Potentials

### v1

@@ 5 Summary
In this work, we train a machine learning interatomic potential (MLIP) foundation model using large-scale molecule relaxation data. By curating a dataset comprising 3.5 million small molecules and 300 million snapshots with energy and force labels computed at multiple levels of quantum accuracy, we enable the development of MLIP foundation models that can be used to efficiently obtain low-energy 3D structures through geometry optimization for downstream property predictions. Additionally, we introduce geometry fine-tuning as a strategy to mitigate errors and biases introduced during structure relaxation, enhancing the downstream property prediction performance using relaxed 3D geometries. Furthermore, we demonstrate that the MLIP foundation model can be directly fine-tuned for molecular property prediction tasks, further extending its applicability.
@@ Acknowledgments
SJ acknowledges support from ARPA-H under grant 1AY1AX000053, National Institutes of Health under grant U01AG070112, and National Science Foundation under grant IIS-2243850. XFQ acknowledges the support from the Center for Reconfigurable Electronic Materials Inspired by Nonlinear Dynamics (reMIND), an Energy Frontier Research Center funded by the U.S. Department of Energy, Basic Energy Sciences, under Award Number DE-SC0023353. We acknowledge the support of Lambda, Inc. and NVIDIA for providing the computational resources for this project.
Appendix A Appendices
A.1 Benchmarked Methods on PubChemQCR-S
@@
SchNet [Schütt et al., 2018] is a continuous-filter convolutional network that models local atomic correlations using learned filter-generating networks. FAENet [Duval et al., 2023] introduces frame averaging to enforce symmetry compliance in molecular systems, allowing geometric information to be processed without explicit symmetry-preserving architectural constraints. NequIP [Batzner et al., 2022] and Equiformer [Liao and Smidt, 2022] are equivariant neural networks that model interatomic interactions while preserving geometric symmetries. They maintain various types of geometric features during message passing and capture feature interactions using the Clebsch–Gordan tensor product. Equiformer further incorporates attention mechanisms into the equivariant message passing process. SevenNet [Park et al., 2024] is an equivariant model that builds upon the NequIP architecture by introducing a scalable parallelization scheme designed for spatial decomposition in large-scale molecular dynamics (MD) simulations. Allegro [Musaelian et al., 2023] is a highly efficient, local equivariant model tailored for scalable, high-accuracy simulations. It models many-body interactions through a series of tensor products of learned equivariant representations. PaiNN [Schütt et al., 2021] extends SchNet by introducing equivariant representations, enabling the model to capture directional dependencies while maintaining computational efficiency. MACE [Batatia et al., 2022] and PACE [Xu et al., 2024] are other equivariant frameworks that capture many-body interactions using symmetry-aware neural architectures.
A.2 Metrics for Geometry Optimization
@@
A.3 Downstream Datasets Details
Molecule3D. The Molecule3D dataset [Xu et al., 2021] is curated from the PubChemQC dataset, which contains nearly 4 million organic small molecules. Each molecule is associated with a ground-state 3D geometry derived from DFT calculations, along with corresponding quantum properties. In our experiments, we focus on predicting the HOMO-LUMO gap, a key quantum property of molecular electronic structure. Molecule3D provides two standard data splits: a random split, where training, validation, and test sets are sampled from the same distribution, and a scaffold split, which introduces a distribution shift between training and test sets to evaluate model generalization. Since Molecule3D and our curated dataset are derived from the same source, we remove any overlapping molecules from the Molecule3D test set to prevent data leakage.
$\nabla^{2}$. $\nabla^{2}$ [Khrabrov et al., 2024] is a recently introduced large-scale benchmark that includes DFT-level ($\omega$B97X-D/def2-SVP) calculations of energies, forces, molecular properties, and Hamiltonian matrices. For training, we use the large split comprising 99,018 molecules and 500,552 conformations. The structure test split includes 176,001 molecules.
A.4 Network and Training Details
For the MLIP foundation model, we use PaiNN with a hidden dimension of 128, 128 Gaussian components in the radial basis function, 4 layers, and a cutoff distance of 4.5Å. We adopt the PaiNN implementation from the Open Catalyst Project GitHub repository [Chanussot et al., 2021].
For pre-training the MLIP foundation model on PubChemQCR, we use a learning rate of 1e-3 and a batch size of 64. Optimization is performed using Adam [Kingma and Ba, 2014] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and learning rate scheduling via ReduceLROnPlateau with a patience of 2 epochs. Training is conducted for 9 epochs on four NVIDIA H100 GPUs.
@@ A.4 Network and Training Details
A.5 Additional Results
A.5.1 Experimental Results of Molecule3D Full Dataset
Table 8 presents the fine-tuning results on the full Molecule3D dataset. With more training data, all methods show improved performance, and the MLIP foundation model continues to achieve the best results, demonstrating the effectiveness of the pre-trained foundation model.
Table 8: Results of HOMO-LUMO gap prediction on two splits of the Molecule3D full dataset. Force2Prop w/ PaiNN denotes the MLIP foundation model pre-trained on the PubChemQCR dataset. Best results are shown in bold, and second-best results are underlined.
@@ A.5.1 Experimental Results of Molecule3D Full Dataset
A.5.2 Different Pre-Training Strategies
In this work, we adopt supervised pre-training to train the foundation model for explicit energy and force prediction. Previously, due to the lack of large-scale relaxation datasets with both energy and force labels, prior methods relied on self-supervised pre-training. Among them, denoising pre-training has been the most effective, and we compare our approach against this strategy. Specifically, we consider two denoising methods: coordinate denoising [Zaidi et al., 2022] and the more recent DeNS method [Liao et al., 2024], which is designed to generalize to non-equilibrium structures. We then fine-tune the pre-trained models on the Molecule3D dataset for molecular property prediction. Results in Table 9 show that supervised pre-training outperforms denoising-based methods on downstream tasks, as it enables the model to explicitly learn atomic interactions.
@@ A.5.2 Different Pre-Training Strategies
A.5.3 Different Training Hyperparameters in Geometry Fine-Tuning
As described in Section 2.4, we perform geometry fine-tuning using multi-task learning by incorporating an additional geometry alignment task. During training, noise is added to the ground truth conformers to improve generalization. We evaluate the impact of different noise scales, with results shown in Table 10. The results indicate that performance is sensitive to the noise level—both overly small and overly large noise can degrade performance. We also investigate the effect of varying the loss weight for the geometry alignment task. As shown in Table 11, the performance is relatively robust to the choice of loss weight.
@@ A.5.3 Different Training Hyperparameters in Geometry Fine-Tuning
A.5.4 Dataset with Only 2D Graphs
In Section 2.4, the task assumes access to ground-truth 3D conformers during training but not at test time. A more challenging setting arises when neither the training nor the test set includes ground truth 3D geometries—only 2D molecular graphs are available. For this experiment, we predict the HOMO-LUMO gap on the QM9 dataset using only 2D graphs. Initial 3D conformers are generated via RDKit and relaxed using our foundation model; the downstream predictor is then trained from scratch on these relaxed geometries. As shown in Table 12, a significant performance gap remains between using ground-truth versus relaxed conformers. This is partly because the foundation model was trained only on geometries near the energy minimum, making relaxation from RDKit conformers an out-of-distribution task. Also, the geometry optimization capability needs further improvement when it is used to generate stable conformers for downstream property prediction tasks. If the optimized geometries are imperfect, correcting the relaxation bias becomes highly challenging without ground-truth geometries available during training. Although our current foundation model does not perform well in this setting, we highlight this scenario as an important future direction, aiming to eliminate reliance on DFT-derived geometries entirely.

### v2

@@ 5 Summary
In this work, we train a machine learning interatomic potential (MLIP) pre-trained model using large-scale molecule relaxation data. By curating a dataset comprising 3.5 million small molecules and 300 million snapshots with energy and force labels computed at multiple levels of quantum accuracy, we enable the development of MLIP pre-trained models that can be used to efficiently obtain low-energy 3D structures through geometry optimization for downstream property predictions. Additionally, we introduce geometry fine-tuning as a strategy to mitigate errors and biases introduced during structure relaxation, enhancing the downstream property prediction performance using relaxed 3D geometries. Furthermore, we demonstrate that the MLIP pre-trained model can be directly fine-tuned for molecular property prediction tasks, further extending its applicability.
6 Broader Impact Statement
DFT-based geometry optimization can produce highly accurate molecular structures for property prediction, but it is computationally expensive. Our MLIP model, trained on the curated relaxation dataset, demonstrates strong potential for efficiently generating approximate geometries and can improve downstream property prediction to a certain extent. However, we emphasize that the geometries produced by the MLIP model are not yet comparable to those obtained via DFT, and caution should be exercised when applying them in critical or high-stakes scenarios. To support further research, we will release both the curated dataset and the trained model to facilitate continued development of MLIP methods.
@@ Acknowledgments
SJ acknowledges support from SES AI Corporation, ARPA-H under grant 1AY1AX000053, National Institutes of Health under grant U01AG070112, and National Science Foundation under grant IIS-2243850. XFQ acknowledges support from SES AI Corporation, the Center for Reconfigurable Electronic Materials Inspired by Nonlinear Dynamics (reMIND), an Energy Frontier Research Center funded by the U.S. Department of Energy, Basic Energy Sciences, under Award Number DE-SC0023353. We acknowledge the support of Lambda, Inc. and NVIDIA for providing the computational resources for this project.
Appendix A More Details about the Large-Scale DFT Relaxation Dataset
In this section, we introduce more details about our curated PubChemQCR dataset, covering data generation, curation process, and data statistics. Data are publicly available at: https://huggingface.co/datasets/divelab/PubChemQCR.
A.1 Dataset Generation
The raw trajectory data are sourced from the PubChemQC database (Nakata & Shimazaki, 2017). The geometry optimization of each molecule follows a structured protocol. First, initial 3D molecular structures are generated from the InChI representation using OpenBabel, providing the starting point for subsequent quantum calculations. The first optimization is performed using the PM3 semi-empirical method. Next, the relaxed structures undergo further refinement using the Hartree-Fock method with the STO-6G basis set. Finally, the 3D structures are fully optimized using the B3LYP functional with the 6-31G* basis set. PM3 and Hartree-Fock optimizations are performed using the GAMESS software. The DFT optimization consists of three substeps: first, Firefly or SMASH is used for a faster but slightly less accurate optimization. Then, GAMESS is employed for more precise geometry optimization, followed by a final validation step to ensure molecules are indeed optimized.
A.2 Dataset Curation
The raw trajectory data is 7TB in size and not immediately suitable for machine learning. To make it more accessible, we parsed all raw log files to extract the energies and forces at each snapshot, along with the atomic numbers and coordinates of each atom.
We save the parsed trajectories in six Lightning Memory-Mapped Database (LMDB) files, leveraging its efficient key-value storage and fast data retrieval. Each trajectory is saved as a key-value pair, where the key is the PubChem CID, a unique identifier for chemical compounds in the PubChem database, and the value is a dictionary containing the parsed trajectory, with keys corresponding to the names of each optimization stage, i.e. pm3, hf, DFT_1st, and DFT_2nd. By storing snapshots from each optimization stage separately, we provide more flexibility in selecting specific trajectory segments for training. The curated dataset is 400GB in size, significantly smaller than the raw data. This size reduction, along with the structured data format, enhances accessibility and usability for the research community, making it easier to develop models using the dataset.
A.3 Dataset Statistics
The full dataset contains 3,471,000 trajectories and a total of 298,751,667 molecular snapshots, including 163,015,359 snapshots from PM3, 19,274,130 snapshots from Hartree-Fock, 105,494,671 snapshots from the first substage of DFT calculations and 10,967,507 snapshots from the second substage of DFT calculations. On average, each molecule consists of 29 atoms, including 14 heavy atoms, and each molecular trajectory contains approximately 47 PM3 snapshots, 6 Hartree-Fock snapshots, 31 DFT first substage snapshots, and 3 DFT second substage snapshots. This dataset covers 25 chemical elements.
For training efficiency, we also curated a smaller subset for model benchmarking, named PubChemQCR-S, which contains 40,979 trajectories and 1,504,431 molecular snapshots from the first substage of DFT calculations.
Appendix B Benchmarked Methods on PubChemQCR-S
@@
SchNet (Schütt et al., 2018) is a continuous-filter convolutional network that models local atomic correlations using learned filter-generating networks. FAENet (Duval et al., 2023) introduces frame averaging to enforce symmetry compliance in molecular systems, allowing geometric information to be processed without explicit symmetry-preserving architectural constraints. NequIP (Batzner et al., 2022) and Equiformer (Liao & Smidt, 2022) are equivariant neural networks that model interatomic interactions while preserving geometric symmetries. They maintain various types of geometric features during message passing and capture feature interactions using the Clebsch–Gordan tensor product. Equiformer further incorporates attention mechanisms into the equivariant message passing process. SevenNet (Park et al., 2024) is an equivariant model that builds upon the NequIP architecture by introducing a scalable parallelization scheme designed for spatial decomposition in large-scale molecular dynamics (MD) simulations. Allegro (Musaelian et al., 2023) is a highly efficient, local equivariant model tailored for scalable, high-accuracy simulations. It models many-body interactions through a series of tensor products of learned equivariant representations. PaiNN (Schütt et al., 2021) extends SchNet by introducing equivariant representations, enabling the model to capture directional dependencies while maintaining computational efficiency. MACE (Batatia et al., 2022) and PACE (Xu et al., 2024) are other equivariant frameworks that capture many-body interactions using symmetry-aware neural architectures.
Appendix C Training Details of Benchmarked Methods
On the PubChemQCR-S subset, Equiformer uses a separate prediction head to directly predict atomic forces, whereas other methods compute forces as the gradient of the predicted energy. Note that to eliminate the influence of molecular size, we predict the energy per atom rather than the total energy. Additionally, we normalize the energy by subtracting the mean energy during training. To remove the effect of translation, we also center the coordinates by shifting them to have a zero centroid.
Table 8: Model configurations—including the number of layers, hidden dimensions (or maximum irreducible representation channels), and batch sizes—are provided for all baseline models trained on the PubChemQCR-S dataset. These models include SchNet (Schütt et al., 2018), PaiNN (Schütt et al., 2021), MACE (Batatia et al., 2022), Equiformer (Liao & Smidt, 2022), PACE (Xu et al., 2024), FAENet (Duval et al., 2023), NequIP (Batzner et al., 2022), SevenNet (Park et al., 2024), and Allegro (Musaelian et al., 2023).
Models Layers Hidden Dimension Batch Size
SchNet 4 128 128
PaiNN 4 128 32
FAENet 4 128 64
NequIP 5 64 16
SevenNet 5 128 16
MACE 2 128 8
PACE 2 128 8
Allegro 2 128 8
Equiformer 4 128 32
[…]
All experiments on the PubChemQCR-S benchmarks employ a cutoff radius of $4.5\,\text{\AA }$, the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of $5\times 10^{-4}$, and a ReduceLROnPlateau learning rate scheduler with a patience of 2 epochs. Models are trained for up to 100 epochs on the PubChemQCR-S subset using NVIDIA A100-80GB GPUs.
Appendix D Metrics for Geometry Optimization
@@
Appendix E Downstream Datasets Details
Molecule3D. The Molecule3D dataset (Xu et al., 2021) is curated from the PubChemQC dataset, which contains nearly 4 million organic small molecules. Each molecule is associated with a ground-state 3D geometry derived from DFT calculations, along with corresponding quantum properties. In our experiments, we focus on predicting the HOMO-LUMO gap, a key quantum property of molecular electronic structure. Molecule3D provides two standard data splits: a random split, where training, validation, and test sets are sampled from the same distribution, and a scaffold split, which introduces a distribution shift between training and test sets to evaluate model generalization. Since Molecule3D and our curated dataset are derived from the same source, we remove any overlapping molecules from the Molecule3D test set to prevent data leakage.
[…]
Appendix F MLIP Pre-Trained Model and Training Details
For the MLIP pre-trained model, we use PaiNN with a hidden dimension of 128, 128 Gaussian components in the radial basis function, 4 layers, and a cutoff distance of 4.5Å. We adopt the PaiNN implementation from the Open Catalyst Project GitHub repository (Chanussot et al., 2021).
For pre-training the MLIP pre-trained model on PubChemQCR, we use a learning rate of 1e-3 and a batch size of 64. Optimization is performed using Adam (Kingma & Ba, 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and learning rate scheduling via ReduceLROnPlateau with a patience of 2 epochs. Training is conducted for 9 epochs on four NVIDIA H100 GPUs.
@@ Appendix F MLIP Pre-Trained Model and Training Details
Appendix G Additional Results
G.1 Experimental Results of Molecule3D Full Dataset
Table˜9 presents the fine-tuning results on the full Molecule3D dataset. With more training data, all methods show improved performance, and the MLIP pre-trained model continues to achieve the best results, demonstrating the effectiveness of the pre-trained pre-trained model.
Table 9: Results of HOMO-LUMO gap prediction on two splits of the Molecule3D full dataset. Force2Prop w/ PaiNN denotes the MLIP pre-trained model pre-trained on the PubChemQCR dataset. Best results are shown in bold, and second-best results are underlined.
@@ G.1 Experimental Results of Molecule3D Full Dataset
G.2 Different Pre-Training Strategies
In this work, we adopt supervised pre-training to train the pre-trained model for explicit energy and force prediction. Previously, due to the lack of large-scale relaxation datasets with both energy and force labels, prior methods relied on self-supervised pre-training. Among them, denoising pre-training has been the most effective, and we compare our approach against this strategy. Specifically, we consider two denoising methods: coordinate denoising (Zaidi et al., 2022) and the more recent DeNS method (Liao et al., 2024), which is designed to generalize to non-equilibrium structures. We then fine-tune the pre-trained models on the Molecule3D dataset for molecular property prediction. Results in Table˜10 show that supervised pre-training outperforms denoising-based methods on downstream tasks, as it enables the model to explicitly learn atomic interactions.
@@ G.2 Different Pre-Training Strategies
G.3 Different Training Hyperparameters in Geometry Fine-Tuning
As described in Section˜2.4, we perform geometry fine-tuning using multi-task learning by incorporating an additional geometry alignment task. During training, noise is added to the ground truth conformers to improve generalization. We evaluate the impact of different noise scales, with results shown in Table˜11. The results indicate that performance is sensitive to the noise level—both overly small and overly large noise can degrade performance. We also investigate the effect of varying the loss weight for the geometry alignment task. As shown in Table˜12, the performance is relatively robust to the choice of loss weight.
@@ G.3 Different Training Hyperparameters in Geometry Fine-Tuning
G.4 Dataset with Only 2D Graphs
In Section˜2.4, the task assumes access to ground-truth 3D conformers during training but not at test time. A more challenging setting arises when neither the training nor the test set includes ground truth 3D geometries—only 2D molecular graphs are available. For this experiment, we predict the HOMO-LUMO gap on the QM9 dataset using only 2D graphs. Initial 3D conformers are generated via RDKit and relaxed using our pre-trained model; the downstream predictor is then trained from scratch on these relaxed geometries. As shown in Table˜13, a significant performance gap remains between using ground-truth versus relaxed conformers. This is partly because the pre-trained model was trained only on geometries near the energy minimum, making relaxation from RDKit conformers an out-of-distribution task. Also, the geometry optimization capability needs further improvement when it is used to generate stable conformers for downstream property prediction tasks. If the optimized geometries are imperfect, correcting the relaxation bias becomes highly challenging without ground-truth geometries available during training. Although our current pre-trained model does not perform well in this setting, we highlight this scenario as an important future direction, aiming to eliminate reliance on DFT-derived geometries entirely.